LT India Ltd has the following capital structure, which it considers optimal:

Debt  35% 

Equity shares 65%

 Total  100% 

Applicable tax rate for the company is 25%. Risk free rate of return is 6%, average equity

market investment has expected rate of return of 12%. The company’s beta is 1.10. Debt

will bear an interest rate of 9% p.a.

 a. component cost of debt and equity shares assuming that the company does not issue

any additional equity shares.   

b.Weighted Average Cost of Capital (WACC).   



1
Expert's answer
2021-11-19T11:02:58-0500

a. Component cost of debt and equity shares assuming that the company does not issue any additional equity shares are:

Cost of debt is 9%.

Cost of equity is:

"Ke = 0.06 + 1.1*(0.12 - 0.06) = 0.126."

b. Weighted Average Cost of Capital (WACC) is:

"WACC = 0.35*(1 - 0.25)*0.09 + 0.65*0.126 = 0.1055."